OpenStudy (anonymous):

f(x) = 5x^2 - 4x g(x) = -3x^2 - 8 Find: (f + g)(2) Please helpp !!!!

OpenStudy (blurbendy):

substitute 2 in for x in both of those equations, then add the results together

OpenStudy (blurbendy):

actually no,

OpenStudy (blurbendy):

based off this notation, im not sure if they want you to substitute 2 in for the g function first. then take that result and add it to the f function

OpenStudy (anonymous):

thats wat i dont get either, so substitute 2 for g first and the f and then mulitply by 2? @blurbendy

OpenStudy (anonymous):

i just did that and i didnt get one of the choices for the answer, what i dont nderstand is what do i use to substitute for.

OpenStudy (anonymous):

\[f.g(2)\]

OpenStudy (anonymous):

i dont understand @telijahmed

OpenStudy (blurbendy):

what are your answer choices

OpenStudy (anonymous):

-8 , -16 , 24 , 16

OpenStudy (blurbendy):

i would go with -8 then. if you substitute 2 in for both of the functions you get -8 5(2)^2 - 4(2) = 5(4) - 8 = 12 -3(2)^2 - 8 = -3(4) - 8 = -20 12 + (-20) = -8

OpenStudy (anonymous):

thank youu.

OpenStudy (blurbendy):

np